OpenStudy (anonymous):

Im having a problem Integrating dx/(2+x^(1/2)) When i used U substitution u^2=x and 2udu=dx i got 2udu/(2+u) but i realy dont know what to do now?

OpenStudy (anonymous):

Try to substitute: \[y = 2 + \sqrt{x} \\ dy = \frac{1}{2\sqrt{x}}dx = \frac{1}{2(y-2)}dx\]

OpenStudy (anonymous):

yes! I solved it now, thanks! That last step when you changed x^1/2 to (y-2) helpt a lot :D